OpenStudy (anikate):

simplify: 6x^5 -51x^3 -27x

OpenStudy (asnaseer):

ok - first can you spot the common factor between all three terms?

OpenStudy (asnaseer):

you could start by just looking at the constants: 6, 51, 27 what number divides into all of these?

OpenStudy (anikate):

3

OpenStudy (asnaseer):

good, so we can first simplify this as follows:\[6x^5-51x^3-27x=3(2x^5-17x^3-9x)\]do you understand so far?

OpenStudy (anikate):

yes :)

OpenStudy (asnaseer):

gr8! so next try and spot what can be factored out of: \(x^5,x^3,x\) what goes into all of these?

OpenStudy (anikate):

x

OpenStudy (asnaseer):

good - can you write out what you think the simplification will be when we take a single 'x' outside of the braces?

OpenStudy (anikate):

3x(2x^4 -17x -9)

OpenStudy (asnaseer):

almost! the 17x should actually be 17x^2

OpenStudy (anikate):

i no I just forgt

OpenStudy (asnaseer):

:) ok, you should therefore end up with:\[6x^5-51x^3-27x=3(2x^5-17x^3-9x)=3x(2x^4-17x^2-9)\]which can be simplified further

OpenStudy (asnaseer):

you should be able to factorise the terms within the braces

OpenStudy (asnaseer):

think of then as a quadratic equation in \(x^2\)

OpenStudy (asnaseer):

so if you say let \(z=x^2\) then you would have: \(2z^2-17z-9\) Can you factorise this?

OpenStudy (anikate):

yes

OpenStudy (asnaseer):

what do you get?

OpenStudy (anikate):

3x(x-9)(2x+1)

OpenStudy (asnaseer):

great work! but you need to be careful - remember we substituted \(z=x^2\), so what you've really ended up with is:\[3x(z-9)(2z+1)\]and if we replace z with \(x^2\) now, then we get:\[3x(x^2-9)(2x^2+1)\]

OpenStudy (asnaseer):

there is one more simplification that can be made - can you spot it?

OpenStudy (anikate):

the squares?

OpenStudy (asnaseer):

thats it! you're an expert in these I see! :)

OpenStudy (anikate):

haha I'm really not but thx

OpenStudy (anikate):

so what do you do to the squares? just remove them?

OpenStudy (asnaseer):

no - when you said the squares I assumed you meant the \((x^2-9)\) term. this is a difference of two squares since it can be written as: \((x^2-3^2)\)

OpenStudy (asnaseer):

and there is a well know factorisation for this

OpenStudy (asnaseer):

\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (asnaseer):

do you understand?

OpenStudy (anikate):

yup= 3x(x+3)(x-3)(2x+1)

OpenStudy (asnaseer):

remember the LSAT term there should be \((2x^2+1)\)

OpenStudy (asnaseer):

*LAST

OpenStudy (anikate):

how to I factor that one?

OpenStudy (asnaseer):

you cannot - you have now simplified as far as possible.

OpenStudy (anikate):

oh ok

OpenStudy (asnaseer):

so final result should be:\[3x(x+3)(x-3)(2x^2+1)\]

OpenStudy (anikate):

so can you plz write the answer :)

OpenStudy (anikate):

oh ok thx also since u r a moderator do u get paid for this?

OpenStudy (anikate):

intern?

OpenStudy (asnaseer):

no :) I do this as a hobby because I enjoy teaching (and learning) maths

OpenStudy (anikate):

so how did u become a moderator?

OpenStudy (asnaseer):

this has been asked a lot in the OpenStudy Feedback Group - let me find a link to it for you...